At a certain temperature, the K for the decompositon of H2S is 0.759.

H2S(g) <---> H2(g) + S(g)

Initally, only H2S is present at a pressure of 0.103 bar in a closed container. What is the total pressure in the container at equilibrium?
Answer #1

Hasig) Halg) + Sly) at t- 0 0103 bas 0 0 at ta teguy 0.103-4 2 since, given that K= 0.759 and we know that X doesnt change w9) x = -0.759 + Jo.888 - 9.759 fo.942 2 - x = -0.759to.942 a 0.759 -0.942 2 2 x = 8.0916 since total preouse at Equi = Proses